OpenStudy (leozap1):

URGENT HELP PLEASE Meddle & fan to who ever helps me The question is: Choose the correct graph of the function y = 3/4 (2^x) The answer choices are a bunch of graphs that look like a exponential function graph so I solved it out like that and got nothing like any of the graphs please help I have no idea how to solve these please help

OpenStudy (leozap1):

Please explain to me how to solve this and please don't give me the answer

OpenStudy (anonymous):

Can you attach the picture maybe that way we could help more.

OpenStudy (leozap1):

These are the answer choices

OpenStudy (anonymous):

if we put x = 0 in y = 3/4 (2^x) we get y = 3/4 = 0.75 and for x = 1 we get y = 3/2 = 1.5 now if we take a close look to the first graph it suits the above points

OpenStudy (leozap1):

I did each one as I would for a exponital function by doing x = -2, x = -1, x = 0, x =1, x = 2 and then put down the points but none of them match the graph that I made so I think there is a different way to do it

OpenStudy (anonymous):

first of all the the function y = 3/4 (2^x) is an exponential raise so we can omit the last choice. now in the above 3 choices the choice we are left is the first one. i can send you the graph from wolfarmalpha.com

OpenStudy (leozap1):

I have no idea how to do this :c

OpenStudy (anonymous):

you can pick the first choice its correct

OpenStudy (leozap1):

I wanted to know how to do it not the answer X_X

OpenStudy (anonymous):

so as i did it for you, by putting values in x we can approximate the graph.

OpenStudy (leozap1):

I don't understand how to do these types of problems still though

OpenStudy (ankit042):

question is \[\frac{ 3 }{ 4*2^{x}}\] right? do you know how to plot a graph for 2^-x ?

OpenStudy (leozap1):

That's not the question @ankit042

OpenStudy (ankit042):

Then is it \[\frac{ 3 }{ 4 }2^{x} \] try to find value of function for diff value of x example x =2 so f(2) =3

OpenStudy (leozap1):

Do I multiply it out like I do for a exponential function?